OpenStudy (swaugh):

can someone check my answer please What is the slope of a line that passes through the coordinates (1,4) and the coordinates (-1,14) Hint: Think rise over run. -5 -1/5 1/5 5-my answer

OpenStudy (ssuzy):

yes

OpenStudy (ssuzy):

its right

OpenStudy (susanna):

the answer is -5

OpenStudy (anonymous):

you can always check things like this using the point slope form

OpenStudy (anonymous):

or using this:

OpenStudy (susanna):

\[\frac{ 14-4 }{ -1-1 }=\frac{ 10 }{ -2 }=5\]

OpenStudy (susanna):

-5*

OpenStudy (anonymous):